this is a chem. 102 question...
n(P1/P2)=deltaH/R(1/T2-1/T1) and I have to solve for T2
thank you

Answer by Nate(3500)

n(P1/P2)=deltaH/R(1/T2-1/T1)
n(P1/P2)/(deltaH/R)= 1/T2-1/T1
n(P1/P2)/(deltaH/R)+(1/T1)= 1/T2
T2(n(P1/P2)/(deltaH/R)+(1/T1))= 1
T2 = 1/(n(P1/P2)/(deltaH/R)+(1/T1))
You simply need to rearrange the variables!